### v1

# thanks: Corresponding author

Abstract
We investigate signatures of quantum chaos and complexity in the quantum annealing Ising model on random Erdős-Rényi graphs. By tuning the connectivity of the graph, the dynamics can be driven from a localized phase through a chaotic regime to an integrable limit. While this dynamical transition reflects in the spectral characteristics, we pursue a broader suite of quantum chaos indicators, some of which can be measured on near-term quantum devices. We study deep thermalization of a quantum state ensemble obtained from a natural unraveling of the subsystem density matrix as an indicator of chaotic dynamics. This extends the analysis of quantum chaos to the ensemble of quantum states. Furthermore, we analyze the eigenstate and eigenvalue correlations through the partial spectral form factor of subsystems and observe distinct signatures of the onset of chaos and its system size dependence, providing experimentally measurable indicators of the localization-to-chaos transition. As a locality-independent probe, we show that the Krylov complexity of operators is also maximized in the chaotic regime, providing a link between graph topology and information scrambling. Finally, we investigate a quantum analogue of the Mpemba effect, where initially “hotter” states can thermalize anomalously fast, a phenomenon most cleanly observed within the chaotic phase. However, away from the chaotic regime, the system is distinguished by multiple crossings across connectivity in its distance from the thermal ensemble with time. Collectively, this work presents a broad characterization of chaos, providing insight beyond spectral analysis and practical indicators for benchmarking near-term quantum devices.
I Introduction
The study of the quantum Ising model has played a central role in modern quantum many-body physics, with implications spanning diverse themes like optimization [1, 2, 3], quantum phase transitions [4], and glassy physics [5, 6, 7]. These models have found practical applications in quantum optimization through quantum annealing algorithms (QAA), which leverage quantum fluctuations to solve challenging optimization problems [8, 9, 10, 11] in noisy intermediate-scale quantum devices. Modern quantum devices exhibit complex graph topologies that emerge from two complementary origins. First, in the encoding of general optimization problems as Ising Hamiltonians [12, 13, 14], the problem structure dictates the qubit interactions. The second stems from hardware constraints in superconducting quantum processors [15, 16], which impose sparse connectivity patterns that resemble semi-random graphs. In such devices, operator spreading [17, 18, 19] and information scrambling [20, 21, 22, 23, 24] correlate with their performance [25, 26, 27]. Operator spreading and information scrambling have been investigated in regular lattice systems with disorder in the context of localization-delocalization transitions [19, 22, 28, 29, 30, 31], but studies on irregular graph topologies are limited. This, therefore, calls for a deeper understanding of how graph topology affects quantum many-body dynamics and the dynamical phases in such devices.
The QAA Hamiltonian generally encodes an Ising-type cost function, where the relative strength of interactions is tuned with time. The connectivity structure of the underlying network is reflected in the spectral properties and dynamics of the quantum Ising system, which serves as the instantaneous QAA Hamiltonian. While sparse connectivity can lead to localized phases, intermediate connectivity regimes exhibit signatures of quantum chaos [32]. In the chaotic regime, the energy level-statistics follow the predictions of random matrix theory (RMT) [33] and the eigenstate thermalization hypothesis (ETH) [34, 35, 36] predicts the eventual approach to the Gibbs ensemble at thermal equilibrium. In the limit of all-to-all connection (in the absence of longitudinal field), the system becomes an integrable Lipkin-Meshkov-Glick (LMG) model [37] exhibiting Poisson level-statistics and breakdown of quantum chaos.
Recent theoretical and experimental developments have extended our understanding of quantum chaos beyond spectral statistics to encompass higher-order correlations in quantum state ensembles, dynamics of operator spreading, and the emergence of complexity under unitary time evolution. In this work, we present an extensive investigation of these aspects in the quantum annealing Ising model on random Erdős-Rényi graphs. We parameterize the connectivity of the graphs with connectance ($\tilde{M}$), which is the connected edges as a fraction of all possible edges. Here, the disorder arises solely from the connectivity pattern instead of interaction strength or external field.
Conventional studies on quantum thermalization focus on the density matrix and the expectation values of local observables at equilibrium predicted by ETH [34, 35, 36]. The recently introduced projected ensemble (PE) [38, 39] provides a natural unraveling of the density matrix to an ensemble of pure quantum states. This allows analysis at the level of the ensemble of states rather than its first moment alone (which is the density matrix). For systems without any non-trivial conserved charges, the PE constructed from the time-evolved state is expected to converge to the Haar ensemble. We obtain the trace distance of the PE from the Haar ensemble as a function of the the system size and $\tilde{M}$ and show that the trace distance shows the fastest decay with the system size at intermediate $\tilde{M}$, indicating the convergence of PE to the Haar ensemble in the chaotic regime. At both the sparse and dense limits ($\tilde{M}\to 0,1$), the trace distance shows a very slow decay with system size. Next, we focus on the partial spectral form factor (pSFF), which goes beyond the spectral form factor (SFF) to capture the correlation among both eigenstates and eigenvalues in the subsystem. PE and pSFF have been demonstrated on multiple platforms [40, 41, 42, 39, 43], motivating the indicators investigated here from an experimental perspective. Thus, they provide a practical benchmarking protocol for quantum annealing devices. To study operator scrambling, we calculate the Krylov complexity (KC) and its late-time saturation with $\tilde{M}$. In the Krylov basis for operators, the initial local operator spreads through a set of Krylov basis states, much like the spread of a particle on a tight-binding model with nearest-neighbor hopping. We observe that the Lanczos coefficients, which are the hopping parameters in the tight-binding notation, display enhanced fluctuations near integrable limits. The KC saturation values are significantly larger in the chaotic regime compared to localized phases. Finally, we investigate the relaxation of simple product states and uncover instances of anomalous thermalization. We observe instances where “hotter” initial states relax more rapidly than “cooler” ones, which is a defining feature of the quantum Mpemba effect (QME). In the chaotic regime, the anomalous relaxation manifests as a smooth, unambiguous decay to equilibrium. This contrasts sharply with the near-integrable and localized limits, where persistent oscillations in the trace distance obscure the relaxation trend. Taken together, this suite of probes provides a detailed picture of the intricate relationship between graph topology and the signatures of quantum thermalization.
This article is organized as follows. Section II provides a brief introduction to the model and the known spectral characteristics. We utilize the projected ensemble protocol to examine deep thermalization through higher-order moments of quantum state ensembles in Sec. III. In Sec. IV, we study the partial spectral form factor (pSFF) of subsystems to identify characteristic signatures of chaotic behavior and its implications in the context of experiments. We also explore the effect of the choice of subsystem on pSFF in this section. Section V presents results on operator delocalization through Krylov complexity. Finally, in Sec. VI we study the relaxation dynamics of initial product states, including the demonstration of a QME. Finally, we conclude in Sec. VII, remarking on the limitations and possible extensions of our study.
II Ising Model on random graph
Figure 1: A single realization of Erdos-Renyi graph with $L=8$ nodes and connectance, $\tilde{M}=0.6$ which implies number of connected edges, $M=17$. Interacting pairs are represented by connected edges.
We consider Ising model as an instantaneous Hamiltonian for the quantum annealing algorithm, which can be expressed as follows,
$H=\frac{J}{M}H_{p}+\frac{g}{L}H_{x}+\frac{h}{L}H_{z}\ ,$ (1)
$H_{x}=-\sum_{i}X_{i},\quad H_{z}=-\sum_{i}Z_{i},\quad H_{p}=-\sum_{i,j}A_{ij}Z_{i}Z_{j}\ ~.$
$\{X_{i},Y_{i},Z_{i}\}$ are Pauli matrices acting on site $i$ and $g,h$ are the strengths of the transverse and longitudinal field, respectively. $A_{ij}$ is the adjacency matrix for the underlying graph where spins are placed on nodes. $A_{ij}=1(0)$ if there is an edge (no edge) between nodes $i$ and $j$. Here, edges between two nodes represent interaction between the spins placed on them. We take $J=1$ throughout our discussion. The adjacency matrices are chosen to represent random Erdős-Rényi graphs [44] $\mathcal{G}(L,M)$ where $L$ and $M$ denote the number of nodes and edges, respectively. $M$ can be obtained as, $M=\frac{1}{2}\sum_{i,j}A_{ij}$. The denominators $L,M$ in the Hamiltonian ensure that the bandwidth of the energy spectrum scales as $\approx O(JL^{0}M^{0})$. For $g=0$, there is a global $U(1)$ symmetry. For generic $g$, at $h=0$, the model has a global parity symmetry given by, $\mathcal{P}=\prod_{i=1}^{L}X_{i}$. All edges have equal strength ($J$) and graph-to-graph variation arises entirely from the adjacency matrix. For $L$ sites, the total number of possible edges is $\binom{L}{2}$. When $M$ edges are selected, there are $\binom{N_{\text{max}}}{M}$ possible random graphs where $N_{\text{max}}=\binom{L}{2}$. We sample these with equal probability and define the connectance $\tilde{M}=M/N_{\text{max}}$ to characterize the lattice graphs in subsequent discussion. $\tilde{M}\in[0,1]$ where $\tilde{M}=0,1$ corresponds to non-interacting and integrable LMG [37] limit respectively (the presence of small longitudinal field causes deviation in spectral characteristics from the LMG model in finite systems). Figure 1 illustrates a single realization of such a lattice. We take $J=1,g=1$ and $h$ to be a small positive value, thus working with a ferromagnetic Hamiltonian. The dynamical properties studied in this work characterize the bulk of the spectrum, which is expected to be independent of the ferromagnetic or anti-ferromagnetic nature of the ground state. Note that the QAA usually targets the edge of the spectrum (ground state).
The onset of chaos with $\tilde{M}$ in this model (with $h=0$) has been studied [32] through spectral analysis. The chaotic behavior sets in for intermediate $\tilde{M}$ and the spectrum shows localization for extreme $\tilde{M}$. For finite $L$, the system appears integrable at $M\to 0,1$. Through spectral analysis and perturbative arguments, it was shown that the distribution of the level-spacing in the unfolded spectrum converges to the Wigner-Dyson(WD) distribution at large $L$ for $\tilde{M}\lesssim 1$ and $\tilde{M}\gtrsim 0$ for intermediate $g(0.5J<g<4J)$. At very large $g$, the system is effectively non-interacting with the eigenstates given by simple product states; thus, the level-spacing ratio starts deviating from WD. At extreme $\tilde{M}$, the level spacing ratio resembles a Poisson distribution, signaling the breakdown of quantum chaos.
We choose $g=1$ throughout the discussion to retain the possibility of chaotic behavior in the system sizes studied ($6\leq L\leq 14$). We will take $h=0.1$ to break the parity symmetry unless mentioned otherwise. Having established the Hamiltonian, we first investigate its thermalization properties at the subsystem level using the projected ensemble and the partial spectral form factor. We then turn to the system’s dynamical response, analyzing the relaxation of both operators and states across the range of connectance, $\tilde{M}$.
III Projected ensemble and emergence of deep thermalization with $\tilde{M}$
At the level of the density matrix, the ETH predicts that subsystems of unitarily evolving isolated systems relax to Gibbs ensembles determined by the initial state. This relaxation emerges from unitary evolution scrambling initially localized information across the entire system, with the subsystem’s complement serving as an effective bath whose degrees of freedom are traced out to obtain the reduced density matrix. However, modern quantum simulator platforms can precisely track all degrees of freedom of the system simultaneously [45, 46, 47], thus making the distinction between the system and bath artificial. Access to the state of the bath allows us to obtain conditional states in the subsystem conditioned on the measurement outcome of the bath. The ensemble of these conditional states is denoted as the Projected Ensemble, which can be considered as a natural unraveling of the reduced density matrix of the subsystem. Unlike the density matrix, which captures only the first moment of the state ensemble, the PE encodes the full distribution of quantum states in Hilbert space, providing access to higher moments and fluctuations of observables beyond their expectation values. This perspective naturally extends the question of quantum thermalization beyond local observables to whether the underlying distribution of quantum states also approaches a universal form. This phenomenon of relaxation of the quantum state distribution, dubbed “deep thermalization”, generalizes the notion of thermalization beyond the density matrix and expectation value of local observables.
To characterize deep thermalization, it is essential to identify the universal ensemble toward which the PE is expected to converge. In the absence of any conservation law, the Haar ensemble [48, 49] is the maximal entropic ensemble of pure quantum states in Hilbert space. For a system initialized in an effective infinite temperature state and evolved under a chaotic Hamiltonian, the subsystem’s reduced density matrix asymptotically approaches the identity, while the moments of the PE simultaneously approach those of the Haar ensemble [38, 39]. The proximity of the PE to the Haar ensemble serves as a measure of deep thermalization. This can be quantified in terms of the distance between the corresponding moments of the two ensembles. Given an ensemble of quantum states $\mathcal{E}=\{P(\psi),\psi\}$, their $k$-th moment is defined as $\rho_{\mathcal{E}}=\sum_{\psi}P(\psi)(|\psi\rangle\langle\psi|)^{\otimes k}$ where $P(\psi)$ is the probability of state $\psi$ in $\mathcal{E}$. The onset of deep thermalization has been explored in various settings including random circuits, generic Hamiltonian system and the systems where ETH breaks down [50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. In the presence of conservation laws or constraints, the asymptotic projected ensemble reached at $L,t\to\infty$ gets modified to the Scrooge ensemble, which is determined by the conserved charges in the initial state and measurement direction [56, 52].
Here, we briefly outline the protocol to construct PE. Consider a system of $L$ sites partitioned into $A$ and $B$ subsystems with $L=L_{A}+L_{B}$. The entire system is described by a pure state, $\psi_{AB}\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ at all times. We repeat a set of projective measurements restricted to the subsystem $B$ on copies of $\psi_{AB}$ and collect the corresponding pure state in $A$ conditioned on the measurement outcome in $B$, $\psi_{A}(b_{i})$
$\psi_{A,b_{i}}=\sum_{b_{i}}\frac{1}{\sqrt{p(b_{i})}}\Pi_{b_{i}}\psi_{AB}=\psi_{A}(b_{i})\otimes|b_{i}\rangle,$ (2)
where $\Pi_{b_{i}}=\mathbb{I}\otimes|b_{i}\rangle\langle b_{i}|$ and $p(b_{i})=\langle\psi_{AB}|\Pi_{b_{i}}|\psi_{AB}\rangle$. $\psi_{A}(b_{i})$ is the conditional pure state obtained in subsystem $A$ provided that the measurement result in $B$ is $b_{i}$. The PE is constructed as, $\mathcal{E}_{\mathrm{PE}}=\{p(b_{i}),\psi_{A}(b_{i})\}$. We can characterize the PE through its $k$-th moment as, $\rho^{(k)}_{\mathrm{PE}}=\sum_{b_{i}}p(b_{i})(|\psi_{A}(b_{i})\rangle\langle\psi_{A}(b_{i})|)^{\otimes k}$. Any moment of observables can be obtained with access to $\mathcal{E}_{\mathrm{PE}}$ as $\langle\hat{O}^{\otimes k}\rangle_{\mathrm{PE}}=\mathrm{Tr}[\rho^{(k)}_{\mathrm{PE}}\hat{O}^{\otimes k}]$. The reduced density matrix of the subsystem (and the expectation values of local observables obtained from it) corresponds to $\rho^{(1)}$, which is described by the ETH for chaotic systems at equilibrium. To compare two state ensembles, we define the trace distance between the $k$-th moments of two ensembles $\mathcal{E}$ and $\mathcal{E^{\prime}}$ as,
$\Delta^{(k)}(\mathcal{E},\mathcal{E^{\prime}})=\frac{1}{2}\mathrm{Tr}\sqrt{(\rho^{(k)}_{\mathcal{E}}-\rho^{(k)}_{\mathcal{E^{\prime}}})(\rho^{(k)}_{\mathcal{E}}-\rho^{(k)}_{\mathcal{E^{\prime}}})^{\dagger}}$ (3)
Overbar denotes quantities averaged over multiple graph realizations. Quantum information is maximally scrambled by chaotic evolution, and we consider that the initial state corresponds to an effective infinite temperature state. In that case, the projected ensemble should approach the Haar ensemble[38, 52], the uniform distribution over all pure states in Hilbert space i.e. achieving maximum entropy in the quantum state distribution. For a system $A$, the Haar distribution as a measure on normalized quantum states $\psi$ can be written as,
[…]
where $d_{A}$ is the dimension of the Hilbert space of subsystem $A$. We consider the proximity of the projected ensemble and Haar ensemble as a metric for chaos and thermalization in this section.
We take initial states in the direct product form as given below,
$|\psi\rangle=\frac{1}{\sqrt{2^{L}}}\otimes_{i=0}^{L}(|0\rangle+\imath|1\rangle)$ (5)
All projective measurements in $B$ are performed in a Pauli-$Z$ basis. A corresponding inverse temperature $\beta$ can be associated with the pure state given as, $E(\psi)=\langle\psi|H|\psi\rangle\propto\mathrm{Tr}(He^{-\beta H)}$. For a traceless Hamiltonian (like in Eq. 1), the mean of the energy spectrum is 0 i.e., zero-energy states correspond to the middle of the eigen spectrum with $\beta=\infty$. The initial state taken in Eq. 5 is such a zero-energy state which is time-evolved with the Hamiltonian in Eq. 1 with h=0.1,g=1h=0.1,\ g=1. Since there are no other conserved charges, the reduced density matrix of a subsystem corresponds to an infinite-temperature density matrix given as the identity. Note that the choice of other measurement bases does not qualitatively affect the convergence of PE in this setup due to the absence of conserved charges in the chaotic regime [56, 52].
Figure 2: (a)-(b) Evolution of $\overline{\Delta^{(1)}}$ with time for $\tilde{M}=0.2$ and $0.4$ respectively. (c)-(d) Evolution of $\overline{\Delta^{(2)}}$ with time for $\tilde{M}=0.2$ and $0.4$ respectively. The overbar suggests quantity averaged over graph realizations. All data represented here are averaged over $500$ graphs. Power-law fits for $L=14$ are shown with black dashed lines, and the corresponding exponents are noted in the plot. Vertical black dotted lines indicate the region where the fitting was done to obtain the power-law exponents. $L_{A}=2$. The parameters for Hamiltonian (Eq. 1) are taken as $g=1.0,h=0.1$.
First, we analyze the convergence of PE to Haar ensemble with time. Figure 2 shows the evolution of trace distance from Haar ensemble with time for the first $2$ moments of PE for multiple $\tilde{M}$. We take $L_{A}=2$ for the subsequent analysis in this section. $\overline{\Delta^{(1)}},\overline{\Delta^{(2)}}$ decays with time for all cases, ultimately saturating to a system size dependent asymptotic value $\overline{\Delta^{(k)}}(t\to\infty)$. The decay with time shows a power-law scaling with the exponent dependent on $\tilde{M}$. The convergence to Haar ensemble is faster in the chaotic regime ($\tilde{M}=0.4$), where it decays with time as $\propto t^{-1.3}$. The exponent decreases as $\tilde{M}\to 0,1$. The power-law exponent appears same for all the moments (upto the error in their estimation). This indicates that the system thermalizes at the same rate for all the moments.
Figure 3: (a)-(c) Asymptotic($t=10^{9}J$) trace distance of first 3 moments of PE from Haar ensemble as a function of $L$. The asymptotic trace distance gets suppressed with increasing $L$ for all $\tilde{M}$. Error bars represent the standard error of mean. (d)-(f) The same data plotted to show the the scaling of $\overline{\Delta^{(k)}}(t\to\infty)$ with $L$ across all $\tilde{M}$ for first $3$ moments of PE. Each data point shown is an average over 500 graph realizations, calculated for a subsystem of size $L_{A}=2$ with Hamiltonian (Eq. 1) parameters $g=1.0$ and $h=0.1$.
Next, we focus on the trend of asymptotic(late-time) trace distance $\overline{\Delta^{(k)}}(t\to\infty)$ with $\tilde{M}$. We compare the asymptotic trace distance between the first three moments of PE and the corresponding moments of the Haar ensemble in Fig. 3(a)-(c) as a function of $\tilde{M}$ for multiple $L$. The asymptotic trace distance is large at extreme values of $\tilde{M}$ (where WD spectral statistics breaks down) and goes down at intermediate $\tilde{M}$ where the spectrum is expected to be chaotic. The results are obtained after averaging over $500$ graph realizations for each $\tilde{M}$. The trace distance has very weak dependence on $L$ at $\tilde{M}\to 0,1$ i.e., where the system approaches an integrable or localized limit. The $\Delta^{(k)}(t\to\infty)$ shows slight asymmetry with $\tilde{M}$ with a minimum appearing around $\tilde{M}\approx 0.5$. Finally, in Fig. 3(d)-(f), we present the scaling of $\overline{\Delta^{(k)}}(t\to\infty)$ with $L$. The asymptotic trace distance from the Haar ensemble decays exponentially with $L$ for intermediate $\tilde{M}$. This supports the convergence of PE to the Haar ensemble at the thermodynamic limit, consistent with the deep thermalization in generic chaotic models at $L\to\infty$ [38, 39]. The decay exponent with $L$ depends on $\tilde{M}$. The trace distance decays weakly with $L$ for $\tilde{M}\to 0,1$, while it shows the fastest decay with $L$ at intermediate $\tilde{M}$. At the fully connected limit, the trace distance is almost independent of system size. The qualitative behavior is the same for all three moments. The breakdown of deep thermalization $\tilde{M}\to 0,1$, can be attributed to the approximate conserved charges close to the integrable limit, which modifies the asymptotic (in $L,t$) ensemble away from the Haar ensemble. This reflects on $\Delta^{(1)}$ itself as the density matrix gets modified to a GGE from the Identity matrix. Note that the minimum of trace distance for a given $L,L_{A}$ appears slightly left of $M=0.5$. This slight asymmetry in asymptotic trace distance as a function of $\tilde{M}$ can be an artifact of small system sizes studied. Similar asymmetry has been observed in level spacing ratio in Ref. [32, 24]).
We have presented the data at the parity-broken ($h=0.1$) limit in this section. When parity symmetry is preserved, measurement outcomes should be post-selected to remain within a given parity sector [55, 51].
IV Dependence of partial SFF on $\tilde{M}$
Having characterized the emergence of chaos through the properties of the state ensemble, we now turn to a complementary diagnostic: the correlations within the energy spectrum itself. To this end, we investigate the partial spectral form factor (pSFF), which generalizes the SFF to subsystems and serves as a probe of spectral correlation, a key signature of quantum chaos. Unlike the full SFF, the pSFF is more readily accessible in experiments. In this section, we utilize the pSFF to study the localization-to-chaos transition as a function of connectivity ($\tilde{M}$) and explore its dependence on both system and subsystem size.
Before discussing the pSFF, it is instructive to review the SFF [60, 61, 18, 62] first. The SFF is a spectral diagnostic of quantum chaos as it captures both short and long-range spectral correlations, providing information beyond that contained in local level-spacing statistics (which captures only information of adjacent level spacings). Formally, SFF is the Fourier transform of the eigenvalue density-density correlation function, defined as follows,
$K(t)=\frac{1}{d^{2}}\sum_{i,j=1}^{d}e^{i(E_{i}-E_{j})t}=\frac{1}{d^{2}}\mathrm{Tr}[T(t)]\mathrm{Tr}[T^{\dagger}(t)]\ ,$ (6)
where $d$ is the Hilbert space dimension of the full system and $E_{i}$ is $i$-th eigenvalue of the Hamiltonian. $t$ is a Fourier variable which can be taken as time. The time evolution unitary operator is represented as, $T(t)=e^{-\imath Ht}$. The prefactor ensures $K(0)=1$. In a chaotic spectra (obeying WD statistics), SFF decays from $1$ at $t=0$ till Thoules time $t_{Th}\approx O(\delta E^{-1})$ where $\delta E$ is the mean level spacing. This results in SFF dropping to below $1/d$ at intermediate time. This is followed by a linear ramp up to saturation to $1/d$ at plateau time (the time at which the plateau in SFF begins), $t_{p}\approx O(2^{L})$. The dip in SFF (referred to as a correlation hole) is caused by level repulsion and spectral rigidity in chaotic systems as predicted by RMT. This ramp-and-plateau structure is a signature of the chaotic spectra. For a spectrum following Poisson statistics, the SFF is expected to gradually reach its asymptotic value after an initial transient period without any correlation hole or ramp. For chaotic systems, the asymptotic SFF is given as $K(t\to\infty)=1/d$ in the absence of any degeneracy in the eigenspectrum. Recently, SFF has been experimentally measured in superconducting qubits using a randomized measurement protocol [40, 41, 63]. Ref [32] analyzes SFF in this model for $h=0$. In the intermediate regime of both $\tilde{M}$ and $g$, the SFF converges with increasing $L$ to the characteristic dip-ramp-plateau structure predicted by RMT, indicating chaotic spectra. Conversely, the chaos is suppressed at the extreme limits of connectivity. In both the sparsely connected ($\tilde{M}\to 0$) and densely connected ($\tilde{M}\to 1$) regimes, the development of the correlation hole is slower with increasing $L$. Similarly, the correlation hole vanishes for both very small and very large $g$, indicating a breakdown of RMT statistics.
However, despite recent experimental demonstrations of SFF measurement on superconducting qubit arrays, a significant practical challenge remains regarding the requirement of global control and site-resolved operations across the entire quantum system. In practice, implementing such fully controlled rotations on all qubits becomes technically demanding and resource-intensive, particularly as system size increases. The partial spectral form factor (pSFF) was recently introduced to address this challenge [40], and unlike the full SFF, it requires controlled operations that scale only with the subsystem size [40, 41]. The pSFF is described in terms of the time-evolution operator on a subsystem. The pSFF can be defined as follows,
$K_{A}(t)$ $=\frac{1}{dd_{A}}\sum_{i,j}e^{\imath(E_{i}-E_{j})t}\mathrm{Tr}_{B}[\rho_{B}(E_{i})\rho_{B}(E_{j})]$
$=\frac{1}{dd_{A}}\mathrm{Tr}_{B}[\mathrm{Tr}_{A}[T(t)]\,\mathrm{Tr}_{A}[T^{{\dagger}}(t)]],$ (7)
where $\rho_{B}(E_{i})=\mathrm{Tr}_{A}\left[|E_{i}\rangle\langle E_{i}|\right]$ is the reduced density matrix of subsystem $B$ for energy eigenstate $|E_{i}\rangle$. pSFF captures correlation of both eigenstates and eigenvalues. Similar to SFF, the normalization factor in Eq.  IV ensures $K_{A}(t=0)=1$.
Figure 4: Time evolution of the pSFF, $\overline{K_{A}(t)}$, for various system sizes ($L$, different colors) and subsystem sizes ($L_{A}$, different line styles). Panels show different connectivities: (a) $\tilde{M}=0.2$, (b) $\tilde{M}=0.4$, (c) $\tilde{M}=0.9$, and (d) $\tilde{M}=1.0$. The solid, dashed, and dotted lines correspond to $L_{A}=1,2,$ and $3$, respectively. The horizontal black lines denote the theoretical asymptotic values for a chaotic system, $1/d_{A}^{2}$. All data are averaged over $500$ graph realizations (denoted by overbar) with Hamiltonian (Eq. 1) parameters $g=1.0$ and $h=0.1$.
We now utilize the pSFF to analyze the Hamiltonian in Eq. 1, setting $g=1$ and $h=0.1$. For the following results, all quantities are averaged over $500$ graph realizations, and the pSFF is calculated on the entire spectrum without an unfolding procedure. We find that, similar to the full SFF, the pSFF universally exhibits an initial polynomial decay. Note that we exclude data from the sparse graphs ($\tilde{M}<0.3$), as the possibility of the subsystem sites belonging to disconnected clusters can cause persistent oscillations in pSFF. We analyze the behavior of pSFF at multiple values of connectivity, $\tilde{M}$, with the results presented in Fig. 4.
First, we describe the intermediate $\tilde{M}$ regime. pSFF shows a dip and a subsequent ramp-and-plateau structure. This correlation hole appears due to correlation in energy spectra, as expected in chaotic systems. The depth of the correlation hole depends on $L,L_{A}$. The depth increases with increasing $L$ and $L_{A}$. The time at the appearance of the correlation hole, $t_{Th}$ increases with $L$ for fixed $L_{A}$. For chaotic systems, RMT predicts [40] (in the absence of degeneracy) $K_{A}(t\to\infty)\approx K(t)+1/d_{A}^{2}=1/d+1/d_{A}^{2}$. For $L_{A}\ll L$, $K_{A}(t\to\infty)\approx 1/d_{A}^{2}$ i,e. asymptotic pSFF shows a subsystem size ($L_{A}$) dependent shift. In Fig. 4, these asymptotic values (for chaotic spectra) are shown using black horizontal lines. For $\tilde{M}=0.4$, Fig. 4(b) shows that the asymptotic $K_{A}$ shows very small fluctuation at late time and approaches $\approx 1/d_{A}^{2}$. Additionally, there is also no significant fluctuation at $t>t_{p}$ in pSFF when no average over graph realization is taken.
In the sparsely connected regime, the pSFF shows qualitatively different behavior from the chaotic case. For $\tilde{M}=0.2$, Fig. 4(a) shows that there is no distinct correlation hole, and the pSFF reaches its asymptotic value after a transient period. The absence of a correlation hole indicates the lack of correlation in the eigenspectrum (consistent with Poisson statistics). There is also a persistent oscillation present at late times after averaging over graph realizations. This can be attributed to quasi-degeneracies in the eigenspectrum (when adjacent level spacing goes below $\delta E$) [32]. For $\tilde{M}$ close to 0, there is a significantly large fluctuation at late time for a single graph realization.
In the densely connected limit approaching all-to-all connectivity, i.e., $\tilde{M}=0.9$, Fig. 4(c) shows a correlation hole in pSFF. In this case, the asymptotic pSFF shows small oscillation around a value larger than $1/d_{A}^{2}$. This deviation from $1/d_{A}^{2}$ is caused by approximate degeneracy in the energy spectrum close to $\tilde{M}=1$. Fig. 4(d) shows the pSFF at $\tilde{M}=1$. A persistent oscillation around at $t>t_{p}$ can be observed. The saturation value at $t>t_{p}$ shows a large deviation from $1/d_{A}^{2}$, hinting at the degeneracy in the spectrum.
At $\tilde{M}=1.0$, i.e., where the model is expected to be almost integrable (in the presence of a small nonzero $h$), we observe a dip in pSFF at an intermediate time. The pSFF eventually saturates to a value larger than the RMT prediction for chaotic systems, suggesting deviation from chaotic spectra.
Figure 5: (Quantitative analysis of the pSFF. (a) Scaling of the plateau time, $\log_{10}t_{p}$, as a function of system size $L$ for several connectance, $\tilde{M}$, with a fixed subsystem size of $L_{A}=2$. (b) The late-time subsystem purity, $\overline{p_{B}}$, as a function of $\tilde{M}$ for a fixed system size $L=12$ and varying subsystem sizes, $L_{A}$. All data are averaged over $500$ graph realizations with Hamiltonian (Eq. 1) parameters $g=1.0$ and $h=0.1$. Here overbar denotes annealed average.
To further quantify the ramp-and-plateau behavior and its dependence on $L$ in the chaotic regime, we study the scaling of the plateau time, $t_{p}$ with $L$ for $L_{A}=2$. This time is determined numerically as the intersection of two linear fits: one to the logarithmic plot of the ramp ($t_{Th}<t<t_{p}$) and another to the late-time plateau ($t>t_{p}$). We observe a clear exponential scaling of $t_{p}$ with system size $L$, as shown in Fig. 5(a), with a slope that does not strongly depend on $\tilde{M}$ within the chaotic regime. The exponential scaling of the plateau time is a feature [40, 41] of the SFF that is inherited by the pSFF. Our data in Fig. 5(a) quantitatively supports this behavior.
Figure 6: Normalized distribution of the asymptotic (late-time) pSFF, $\overline{K_{A}}(t\to\infty)$, computed over all possible two-site subsystems for different system sizes, $L$. Panel (a) shows the distribution in the chaotic regime, $\tilde{M}=0.6$. Panel (b) shows the distribution in the near-integrable regime, $\tilde{M}=0.9$. For all calculations, the subsystem size is fixed at $L_{A}=2$ and the Hamiltonian (Eq. 1) parameters are $g=1.0$ and $h=0.1$.
Another related quantity to pSFF is subsystem purity, defined as $p_{B}=\mathrm{Tr}(\rho_{B}^{2})$ where $\rho_{B}$ is the reduced density matrix of subsystem $B$. This quantity captures the entanglement of subsystem with its complement and is related to subsystem Renyi entropy as $R_{2}=-\log(p_{B})$. The purity can be obtained from pSFF at late time as $p_{B}=\lim_{t_{p}/t\to 0}K_{A}(t)d_{A}$. Figure 5(b) shows $p_{B}$ as a function of $\tilde{M}$ for multiple $L_{A}$ and $L=12$. The purity is minimum at the intermediate $\tilde{M}$ i.e. subsystem $B$ is more mixed (higher entropy) in the chaotic regime compared to the integrable or localized regime. At both $\tilde{M}\to 0,1$, the purity goes up. The purity goes down with increasing $L_{A}$. The higher purity at $\tilde{M}\to 0,1$ is consistent with the expectation of lower entanglement in the localized or integrable phase compared to the chaotic phase. The purity of $B$ decreases (i.e. increase in entanglement of subsystems) with increasing $L_{A}$.
To complete our analysis of the pSFF, we investigate the fluctuation in pSFF due to the choice of subsystem. We compute pSFF for $L_{A}=2$ subsystems over $200$ graph realizations for multiple $L$. In each realization, we choose all possible subsystems consisting of two sites as $A$ and the rest of the sites constitute $B$. Figure 6(a) shows the distribution of asymptotic pSFF values for $L_{A}=2$ and multiple $L$ at $\tilde{M}=0.6$. The variance reduces with increasing $L$. At large $L$, the choice of sites does not affect the asymptotic pSFF in the chaotic regime. The peak also shows a shift towards the expected value (in chaotic spectra) $\approx 1/d_{A}^{2}$ ($0.0625$ for $L_{A}=2$) with increasing $L$. This behavior contrasts sharply with the near-integrable regime ($\tilde{M}=0.9$), where ETH breaks down. Here, Fig. 6(b) shows a broad, multi-peaked distribution for small $L$. The distribution gets sharper with $L$ for fixed $L_{A}$. However, is is significantly broader than the pSFF distribution in the chaotic case for the same $L,L_{A}$.
As the pSFF can be experimentally obtained using randomized measurement, we therefore briefly comment on the number of measurements required to estimate pSFF with sufficient accuracy to capture the physics described in this section. As established in Ref [40], the relative error, $\epsilon$, of the pSFF measurement scales with the number of experimental shots, $N_{m}$, as $\epsilon\propto\sigma_{A}/{K_{A}\sqrt{N_{m}}}$ at each instant of time where $\sigma_{A}$ is the variance in the shadow estimator of pSFF. This implies that the number of measurements needed to reach a desired precision, $N_{m}\propto\sigma_{A}^{2}/K_{A}^{2}$. The number of required measurement scales inversely with the instantaneous value of pSFF. This makes the estimation of the dip in pSFF (i.e around $t=t_{Th}$) challenging for large subsystem in chaotic regime as $K_{A}(t\approx t_{Th})=O(10^{-3})$ in our model. Compared to this, the asymptotic value of pSFF serves as a better distinguishing factor as the magnitude is typically one order of magnitude larger here ($t>t_{p}$). However, obtaining plateau requires long time evolution, making it susceptible to decoherence. The correlation hole appears at an earlier time, though it requires a larger number of measurements.
V Operator delocalization through Krylov complexity
Unitary evolution of a closed quantum system with a generic Hamiltonian results in initial local operators growing into more complex operators with time. While probes like Out-of-Time-Ordered Correlators (OTOCs) [21] can be studied for operator spreading, their reliance on spatial separation is ill-suited for random graphs that lack an inherent local structure. A locality-independent quantification of complexity is the measure of the extent to which an operator spreads in a basis composed of increasingly complicated operators under time evolution [64, 65, 66]. In this section, we analyze the dynamics of operator spreading in the system through Krylov complexity. The time evolution of an operator is generated via application of the propagator constructed with the Liouvillian superoperator, and the resulting KC dynamics provide a diagnostic to quantify chaos and localization.
First, we briefly review KC from the operator viewpoint (An alternative formulation for spread complexity of state has also been proposed [67]). We start with an initial operator $\mathcal{O}(0)$. In Heisenberg picture, the time-evolved operator $\mathcal{O}(t)$ can be expressed as,
$\mathcal{O}(t)=e^{\imath Ht}\mathcal{O}(0)e^{-\imath Ht}=e^{\imath\mathcal{L}t}|\mathcal{O}(0))$ (8)
where Liouvillian, $\mathcal{L}=(\mathbb{I}\otimes H-H^{T}\otimes\mathbb{I})$ and $|\mathcal{O})=\mathrm{vec}(\mathcal{O})\in\mathcal{H}_{d}\otimes\mathcal{H}_{d}$ refers to the vectorized operator defined using,
$\mathcal{O}=\sum_{i,j}\mathcal{O}_{i,j}|i\rangle\langle j|$
$\implies|\mathcal{O})=\text{vec}\left[\sum_{i,j}\mathcal{O}_{i,j}|i\rangle\langle j|\right]=\sum_{i,j}\mathcal{O}_{i,j}|i\rangle|j\rangle.$ (9)
Here, $|i\rangle,|j\rangle$ refers to computational basis states. The action of $\mathcal{L}$ on a generic operator $\mathcal{O}$ is defined as $\mathcal{L}|\mathcal{O})=\mathrm{vec}([H,\mathcal{O}])$. Subsequently, we can define the Krylov subspace ($\mathcal{K}_{\mathcal{O}}$) of $\mathcal{L}$ associated with $\mathcal{O}$ as the minimal subspace where $\mathcal{O}$ is restricted under time-evolution with $\mathcal{L}$ and is spanned by,
$\mathcal{K}_{\mathcal{O}}$ $=\text{span}\{\mathcal{L}^{n}|\mathcal{O})\}_{n=0}^{\infty}=\text{span}\{|\mathcal{O}),|\mathcal{L}\mathcal{O}),|\mathcal{L}^{2}\mathcal{O}),$...\}
$=\text{span}\{\text{vec}(\mathcal{O}),\text{vec}([H,\mathcal{O}]),\text{vec}([H,[H,\mathcal{O}]])$...\}. (10)
The dimension of $\mathcal{K}_{\mathcal{O}}$ is denoted as $K$ and bounded [68] as $1\leq K\leq d^{2}-d+1$. Evidently, the nested commutator operators in $\mathcal{K}_{\mathcal{O}}$ gets more wide (have larger support) with increasing order. We define an inner product on operator space as $(A|B)=\mathrm{Tr}(A^{\dagger}B)$, and work with a complete orthonormal basis $\{|\mathcal{O}_{n})\}_{n=0}^{K-1}$ in $\mathcal{K}_{O}$ obtained using and Lanczos algorithm implemented through succesive full Gram-Schmidt orthogonalization (See Appendix A). Here, $|\mathcal{O}_{0})=\mathrm{vec}(\mathcal{O}(0))$ i.e. the first element of the basis is the initial operator. This orthonormal basis in operator space is denoted as Krylov basis. The Liouvillian takes a tri-diagonal form when expanded in $\{|\mathcal{O}_{n})\}$,
$\mathcal{L}=\begin{pmatrix}0&b_{1}&&&&\\ b_{1}&0&b_{2}&&&\\ &b_{2}&0&\ddots&&\\ &&\ddots&\ddots&b_{K-2}&\\ &&&b_{K-2}&0&b_{K-1}\\ &&&&b_{K-1}&0\end{pmatrix}$ (11)
where diagonal elements are 0 and off-diagonal elements ($\{b_{n}\}$}) form an ordered set of coefficients referred to as Lanczos coefficients. Under time-evolution with $\mathcal{L}$,
$\mathcal{O}(t)=e^{\imath\mathcal{L}t}|\mathcal{O}_{0})=\sum_{n=0}^{K-1}\imath^{n}\phi_{n}|\mathcal{O}_{n})$ (12)
where $\phi_{n}$ are the expansion coefficients of $\mathcal{O}(t)$ in $\mathcal{K}_{O}$. The unitarity of time evolution ensures $\sum_{n=0}^{K-1}|\phi_{n}|^{2}=1$. The time-evolution of $\phi_{n}$ are given by a set of $K$ coupled differential equations,
$\partial_{t}{\phi}_{n}(t)=\phi_{n-1}(t)b_{n}-\phi_{n+1}(t)b_{n+1},$ (13)
where $n\in[0,K-1]$ and $\phi_{n}(0)=\delta_{n,0}$ is taken as initial condition, $b_{0}=b_{K}=0$ is the boundary condition. The form of $\mathcal{L}$ suggests an interpretation of this dynamics analogous to a particle hopping problem on a semi-infinite chain, with $\{|O_{n})\}_{n=0}^{K-1}$ acting as the sites and the hopping amplitudes given by the Lanczos coefficients. To elucidate this, we can represent $\mathcal{L}$ in Krylov basis,
$\mathcal{L}=\sum_{n=0}^{K-1}b_{n}(|\mathcal{O}_{n})(\mathcal{O}_{n-1}|+|\mathcal{O}_{n-1})(\mathcal{O}_{n}|)$ (14)
Figure 7: Analysis of the Lanczos coefficients, $b_{n}$, for an initial operator with support $|\mathcal{O}|=1$. The main panel shows the average coefficients, $\overline{b_{n}}$, as a function of the iteration number $n$ for $\tilde{M}=0.6,0.9$. Error bars represent the standard error of the mean. ( Inset )(a) shows the initial growth of $\overline{b_{n}}$ for the first 20 iterations. (b) shows the distribution of the logarithm of the ratio of consecutive coefficients, $\log(\overline{x})$, where $x_{n}=b_{n}/b_{n+1}$. (c) shows showing the variance of the $\log(\overline{x})$ distribution as a function of $\tilde{M}$. All data are for a system of size $L=6$ with Hamiltonian parameters $g=1.0$ and $h=0.1$, averaged over $500$ graph realizations.
The corresponding evolution of wavefunction representing this particle, $\mathbb{\phi}=\{\ \phi_{0},\imath\phi_{1},\cdots,\imath^{K-1}\phi_{K-1}\}$ can be obtained by solving the set of coupled equations in 13. The KC is defined as,
$\mathcal{C}(t)=\sum_{n=0}^{K-1}n|\phi_{n}(t)|^{2}$ (15)
measuring the particle’s mean position in the Krylov chain. Alternatively, $|\mathcal{O}_{n}\rangle$ involves $n$-th order nested commutators, resulting in Krylov basis elements that have increasingly larger support and more complicated structure with increasing $n$. KC captures the spread of a time-evolved operator in this Krylov basis. We denote the support of an operator $\mathcal{O}$ as $|\mathcal{O}|$.
The Lanczos coefficients $b_{n}$ control how quickly operators spread through the Krylov basis during time evolution. The scaling of $b_{n}$ with $n$ and its fluctuation serve as a signature that can distinguish quantum chaotic dynamics from integrable or localized dynamics [69]. In the following discussion, we take the initial operator in a simple product form.
$\mathcal{O}(0)=\prod_{i=1}^{|\mathcal{O}|}Y_{i}$ (16)
Here, $|\mathcal{O}|$ denotes the size of $\mathcal{O}$ i.e. number of sites where the operator acts non-trivially. Note that we do not have any underlying spatial structure in the lattice; thus, no obvious local structure can be attributed to $\mathcal{O}(0)$ when it has support larger than $1$. In Fig. 7, we show the distribution of $\overline{b_{n}}$ as a function of $n$. $\overline{b_{n}}$ grows at small $n$, saturates at intermediate $n$ and starts descending finally. In a single particle picture, this suggests a hopping problem with site-dependent hopping parameters. The fluctuation in Lanczos amplitude grows as $\tilde{M}\to 1$ i.e., close to the integrable limit. To quantify this fluctuation, we probe the distribution of the ratio of consecutive Lanczos coefficients, $x_{n}=\left(\frac{b_{n}}{b_{n+1}}\right)$. The distribution of $x$ has a mean 0 with the variance increasing as $\tilde{M}\to 1$. In the single particle picture, this indicates the fluctuation in the hopping parameter for adjacent bonds goes up as we approach integrability [70] at extreme $\tilde{M}$. The inset (a) in Fig. 7 shows the growth and subsequent saturation of Lanczos coefficients for small $n$. In the chaotic regime, this saturation is caused due to finite system sizes.
Figure 8: First few Lanczos coefficients and Early-time growth of Krylov complexity. (a) Log-log plot of the average Lanczos coefficients, $\overline{b_{n}}$, versus the iteration number $n$ for several system sizes $L$ at $\tilde{M}=0.6$. Dashed lines show power-law fits, $\overline{b_{n}}\propto n^{\delta}$. All data points are averaged over $200$ graph realizations. (b) The corresponding early-time growth of Krylov complexity, $\overline{\mathcal{C}}(t)$, for the chaotic ($\tilde{M}=0.4$, red) and near-integrable ($\tilde{M}=0.9$, yellow) regimes. The black dashed lines show the linear fit to late-time data. The inset shows a magnified view of the initial non-linear growth. All data are for a single-site initial operator ($|\mathcal{O}|=1$) with Hamiltonian ( Eq. 1) parameters $g=1.0$ and $h=0.1$. For KC, we did a quenched average over $500$ graph realizations.
Figure 9: Evolution and initial growth of Krylov complexity, $\overline{\mathcal{C}}(t)$, for a system with $L=6$. (a)-(c) The time evolution for initial operators with increasing support, $|\mathcal{O}|=1,2,$ and 33 with $\mathcal{O}=\prod\limits^{|\mathcal{O}|}_{i=1}Y_{i}$, respectively, across multiple connectance, $\tilde{M}$. Error bars indicate the standard error of the mean. (d) A magnified view of the initial growth for the same operator sizes shown in (a)-(c). All data are presented after averaging over $500$ graph realizations with Hamiltonian parameters $g=1$ and $h=0.1$ from Eq. 1.
The growth of the Lanczos coefficients, $b_{n}$, directly governs the multi-stage evolution of the Krylov complexity, $\overline{\mathcal{C}}(t)$. For a single-site initial operator, $\mathcal{O}=Y_{1}$, in the chaotic regime ($\tilde{M}=0.6$), we observe that the average coefficients exhibit a power-law growth, $\overline{b_{n}}\propto n^{\delta}$, as shown in Fig. 8(a). With increasing system size $L$, the exponent $\delta$ approaches $1$, indicating the near-linear growth of $\overline{b_{n}}$ that is characteristic of chaotic systems [66, 68]. We observe an even-odd fluctuation about the linear trend in the graph averaged Lanczos coefficients ($\overline{b_{n}}$). Such fluctuations have been observed in quantum field theories [71] as well as Hamiltonian [72, 73] systems earlier. This linear growth of $b_{n}$ is known to cause an initial exponential growth of complexity for times $t\lesssim\log(L)$, a regime physically attributed to the operator scrambling across the system’s real-space support [70, 74]. For the system size studied ($L=6$), the sub-linear growth in $b_{n}$ results in an initial growth of $\mathcal{C}(t)$ that is slower than this ideal exponential behavior. This initial phase is followed by a robust linear growth, indicated by the linear fits (dashed lines) in Fig. 8(b), during which the operator spreads through the higher-order elements of the Krylov basis. Finally, at late times, the finite size of the system halts this spread, causing the complexity to saturate. These qualitative features i.e., initial non-linear growth that transitions into a linear region in KC, and saturation at late time due to finite $L$ are present in both the chaotic and near-integrable regimes, though the initial growth rate of KC slows down near integrability.
The late-time saturation of Krylov complexity, $\overline{\mathcal{C}}(t\rightarrow\infty)$, also serves as a powerful indicator of chaos. The saturation value is expected to be higher for chaotic systems compared to integrable or localized systems, hinting at the emergence of more complicated operators at late times where dynamics are not hindered by constraints or approximate conserved charges. This is evident in Fig. 9(a-c), where the KC saturates to a significantly larger value (in the range of 100014001000-1400) for intermediate $\tilde{M}$, compared to the extreme $\tilde{M}$ where $\overline{\mathcal{C}}(t\rightarrow\infty)\approx 5-200$ for all operator sizes $|\mathcal{O}_{0}|$. The saturation to a value below the theoretical maximum of $K/2$ can be attributed to finite system size and potential degeneracies in the spectrum. We do not observe a clear monotonic dependence of the saturation value on $|\mathcal{O}|$, and these results do not change qualitatively for other initial operators, as shown in Appendix B.
Further, as shown earlier, the initial growth rate of complexity also distinguishes the different dynamical regimes. Figure 9(d) shows that the rate of initial growth is faster in the chaotic regime compared to the near-integrable limits. The growth rate also increases with the support of the initial operator, $|\mathcal{O}_{0}|$. This is physically plausible, as an operator with a larger initial support has already “explored” a larger part of the system’s real space.
Figure 10: The late-time Krylov wavefunction, $|\phi_{n}(t)|^{2}$, shown at $t=2\times 10^{5}$ for several connectance, $\tilde{M}$. The inset provides a magnified view of the occupancy for the first few Krylov basis states at the late time. All data are for a single-site initial operator ($\mathcal{O}=Y_{1}$) on a system of size $L=6$ with Hamiltonian parameters $g=1.0$ and $h=0.1$, averaged over $500$ graph realizations.
The analogy of a particle hopping on the Krylov chain can be made more direct by examining the probability distribution of the particle’s position, given by the Krylov wavefunction, $|\phi_{n}|^{2}$. This is reported in Fig. 10 for multiple $\tilde{M}$ at a late time of $t\approx 2\times 10^{5}$. A significant weight in $|\phi_{n}|^{2}$ at large $n$ indicates that the time-evolved operator, $\mathcal{O}(t)$, has spread to more complex operators in the Krylov basis. We observe that $|\phi_{n}|^{2}$ decays rapidly with $n$ at large $\tilde{M}$, indicating that the operator remains localized near the beginning of the Krylov chain. In contrast, it shows a much slower decay in the chaotic regime, signifying delocalization across the chain. The qualitative trend holds for all operator sizes $|\mathcal{O}|$. Thus, the single-particle picture provides a consistent and intuitive explanation for the operator complexity. The transition to chaos is reflected in the eventual delocalization of the operator across the entire Krylov chain.
VI Anomalous relaxation of initial product states
We now turn to the dynamical relaxation of quantum states under unitary time evolution. In closed quantum many-body systems, a subsystem’s approach to equilibrium is driven by the scrambling of local information, a process whose efficiency was characterized in the previous section via Krylov complexity. In this section, we focus on how certain simple initial states relax under this unitary dynamics. An intuitive expectation for individual states is that the time to reach equilibrium should monotonically increase with the initial state’s “distance” from the equilibrium ensemble. In a classical context, a “hotter” system is farther from ambient temperature compared to a “cooler” system. So, we can expect that a “cooler” state, being “closer” to equilibrium, would relax faster than a “hotter” one. However, it has been found that certain initial conditions can defy this intuition, giving rise to anomalous thermalization behaviors. Broadly, this phenomenon is studied as the Mpemba effect, first reported in water [75], where initially hotter systems freeze faster than cooler ones under a quench.
Recently, the quantum Mpemba effect, a quantum analogue of the counterintuitive classical phenomenon, has been studied. The QME was first proposed for open quantum systems, where a system’s coupling to a thermal bath induces irreversible dynamics [76, 77, 78, 79]. In this setting, the effect is understood to arise from the spectral properties of the system’s dissipative Liouvillian. An initial state can be designed such that its overlap with the slowest eigenmode of the Liouvillian is suppressed, allowing it to equilibrate anomalously fast. This open-system QME was recently experimentally demonstrated [80].
Subsequently, the QME has been extended to isolated systems with global symmetry undergoing unitary evolution, where thermalization is observed at the subsystem level [81, 82, 83, 84, 85]. In such closed systems, the “distance” from equilibrium is typically inferred from the off-diagonal blocks in the symmetry-imposed subsystem reduced density matrix at equilibrium after a quench from an initial state which is not an eigenstate of the global symmetry. Recently, QME has been studied in spin chains and Floquet systems where global symmetries are absent [85] as well as systems with disorder [86].
Figure 11: Distribution of energy of the initial state as a function of $(\theta,\phi)$. $L_{A}=4,L=12$. The initial states are of the form in Eq. 18. The markers on the colorbar indicate the bandwidth of the eigen spectrum.
To investigate the QME in this random graph system, we first establish a quench protocol and a metric to quantify the thermalization. Unlike studies that track the restoration of a global symmetry, our Hamiltonian (Eq. 1) with $h\neq 0$ lacks such a conserved quantity. We therefore adopt a more direct approach [85] to quantify relaxation. Our metric is the trace distance, $\Delta^{(1)}$ (Eq. 3), between the time-evolved reduced density matrix and its corresponding diagonal ensemble. This diagonal ensemble represents the infinite-time equilibrium ensemble ($\rho_{A}(\infty)$) given as,
$\rho_{A}(\infty)=\mathrm{Tr}_{B}(\mathcal{U}\mathrm{diag}(|\mathcal{U}\psi\rangle|^{2})\mathcal{U}^{\dagger}),$ (17)
where $H=\mathcal{U}\tilde{E}U^{\dagger}$ for Hamiltonian $H$ and $\tilde{E}$ is a diagonal matrix containing the eigenvalues of $H$. $B$ indicates the complement of subsystem $A$. To simplify the parameter space of the initial state, we prepare a family of translation-invariant product states $\psi(\theta,\phi)$ in real space as,
$\psi(\theta,\phi)=\otimes_{i=0}^{L}\left(\cos\left(\frac{\theta}{2}\right)|0\rangle+e^{\imath\phi}\sin\left(\frac{\theta}{2}\right)|1\rangle\right).$ (18)
The initial state energy $E(\theta,\phi)$ is a smooth function of $\theta,\phi$ and given as,
$E(\theta,\phi)=-(\cos\theta)^{2}-g\sin\theta\cos\phi-h\cos\theta.$ (19)
$E(\theta,\phi)$ depends only on the coupling parameters of the Hamiltonian but not on the connectance (This is due to the normalization factors in the denominator of Eq. 1). The distribution of initial state energy as a function of $(\theta,\phi)$ is shown in Fig. 11. The red and blue region shows high and low energy states, respectively, with a crossover region near 0 energy where states correspond to infinite temperature. Consider two initial states $\psi(\theta_{1},\phi_{1})$ and $\psi(\theta_{2},\phi_{2})$ such that $\Delta^{(1)}_{\theta_{1},\phi_{1}}(t=0)>\Delta^{(1)}_{\theta_{2},\phi_{2}}(t=0)$ where trace distance is always taken from the respective diagonal ensembles. Existence of QME implies for $t>t_{M}$, $\Delta^{(1)}_{\theta_{1},\phi_{1}}(t)<\Delta^{(1)}_{\theta_{2},\phi_{2}}(t)$, where $t_{M}$ is referred to Mpemba time. Here, $\Delta^{(1)}_{\theta_{1},\phi_{1}}(t=0)$ is the initial “hotter” state due to the initial higher trace distance of the subsystem reduced density matrix from the corresponding equilibrium ensemble.
Our analysis focuses on two distinct classes of initial states(Eq. 18), parametrized by $\theta$ and marked in Fig. 11. The first class, with $\theta=0.5\pi$, spans a wide range of initial energies as $\phi$ is varied. In contrast, the second class, with $\theta=0.95\pi$, consists of nearly-polarized states whose initial energies are largely independent of $\phi$.
Figure 12: Anomalous relaxation in the chaotic regime at $\tilde{M}=0.4$ for a system with $L=12$. (a) The Inverse Participation Ratio (IPR) as a function of the initial state parameters $(\theta,\phi)$. The red and blue vertical lines at $\theta=0.5\pi$ and $\theta=0.95\pi$ indicate the two classes of states studied. (b) Relaxation dynamics, measured by the trace distance $\Delta^{(1)}(t)$ from the diagonal ensemble, for the nearly-polarized states ($\theta=0.95\pi$). (c) The distribution of overlaps, $|\langle E_{n}|\psi\rangle|^{2}$, with energy eigenstates for several initial states from the $\theta=0.5\pi$ class. The inset shows the corresponding IPR for each state. (d) Relaxation dynamics for the $\theta=0.5\pi$ class, showing a clear crossing indicative of the QME. For all plots, the subsystem size is $L_{A}=4$, and all data are averaged over $500$ graph realizations using the Hamiltonian in Eq. 1 with $g=1.0$ and $h=0.1$.
Starting with these two classes of initial states (identified by $\theta$), we time evolve with a specific graph realization of Hamiltonian in Eq. 1 with $h=0.1,g=1$. We first consider $\tilde{M}=0.4$, where the model is expected to be chaotic. Figure 12(b) and (d) shows time evolution of $\overline{\Delta^{(1)}}$ for $\theta=0.5\pi$ and $\theta=0.95\pi$, respectively. All data are presented after averaging over $500$ graph realizations starting with the same set of initial states. We observe a clear crossing in $\Delta^{(1)}$ in Fig. 12(d) where the initial states are taken such that $\theta=0.5\pi$. Such crossing is absent in Fig. 12(b) where $\theta=0.95\pi$ and relaxation to equilibrium looks almost independent of the initial state at the late time.
The observed differences in the late-time relaxation dynamics can be qualitatively understood by considering how the initial states are represented in the energy eigenbasis. $\Delta^{(1)}$ can be expressed in terms of the overlap of the initial state with energy eigenstates as,
$\Delta^{(1)}(t)=\frac{1}{2}\left\|\sum_{n\neq m}c_{n}c_{m}^{*}e^{-i(E_{n}-E_{m})t}\mathrm{Tr}_{\mathrm{B}}(|E_{n}\rangle\langle E_{m}|)\right\|_{1}$ (20)
where $c_{n}=\langle\psi|E_{n}\rangle$. At $t=0$, $\Delta^{(1)}(0)=\frac{1}{2}\left\|\sum_{n\neq m}c_{n}c_{m}^{*}\mathrm{Tr}_{\mathrm{B}}(|E_{n}\rangle\langle E_{m}|)\right\|_{1}$ i.e. the trace distance contains $O(d^{2})$ terms weighted by the overlaps of initial state with eigenstates. As was pointed out in Ref [85], the relaxation rate of states can be related to the number of nonzero coefficients $c_{n}$, quantified by the inverse participation ratio (IPR) given as,
$\mathrm{IPR}(\psi)=\sum_{i=1}^{d}|\langle E_{i}|\psi\rangle|^{4}.$ (21)
If the initial state $\psi$ is a generic state (with energy close to 0, i.e., analogous to a mid-spectrum eigenstate), $c_{n}\approx O(1/\sqrt{d})$ in Eq. 20, corresponding to a smaller IPR.
$|\Delta^{(1)}(t)|^{2}$ is the sum of squared norms of the matrix elements of $(\rho_{A}-\rho_{A}{(\infty}))$. For an initial state $\psi$ which has overlap with a large number of energy eigenstates (i.e., smaller IPR), there are large number of nonzero $c_{n}$s. Each $c_{n}$ is smaller due to normalization of the initial state and, at the same time, there are a larger number of fluctuating phases in the expansion of each matrix elements of $(\rho_{A}-\rho_{A}{(\infty}))$. As a result, the temporal fluctuation in trace distance as the system relaxes are weaker in the case where IPR is smaller. For states with high IPR, a larger number of $c_{n}$s will be close to 0, which results in temporal oscillation in trace distance from equilibrium. This causes a slower equilibration.
In Fig. 12, for the “cooler” initial states, the initial distance from the diagonal ensemble is lower as compared to “hotter” states. Though, we can not prove this, we attempt to provide a plausibility argument. First, we consider the low IPR states. This states have finite expectation value for Pauli observables. The diagonal ensemble constructed with this initial state will be a mixture of highly entangled mid-spectrum eigenstates. However, the diagonal ensemble will have 0 expectation value of such observables (as it is an almost uniform mixture of mid-spectrum eigenstates, which resemble infinite temperature states according to ETH). This suggests that the diagonal ensemble is not a good approximation at the initial time for the direct product state, resulting in a large initial trace distance. Note that, states close to the ground state are being approximated by a diagonal ensemble which, consists of low-lying eigenstates that can have finite local expectation values.
We present IPR as a function of initial state parameters in Fig. 12(a). Interestingly, IPR shows a clear correlation with energy (Fig. 11). The states with energy close to 0 (i.e, they are similar to mid-spectrum states) have a lower IPR (i.e., delocalized in the eigenspectrum). The states away from the middle of the spectrum have higher IPR (i.e., localized in the eigenspectrum). The states considered for time evolution are marked in parameter space with dots. The states with $\theta=0.5\pi$ show a spread in their IPR values. The IPR distribution is consistent with the observation in Fig. 12(d). The states with smaller IPR decay faster in spite of having a larger initial trace distance from their equilibrium density matrix compared to states that have a larger IPR. A smaller IPR signifies a larger number of fluctuating phases in Eq. 20, which explains the faster relaxation. Further, we look at the overlap of the initial states taken in Fig. 12(b) with individual energy eigenstates and present them in Fig. 12(c). The state having energy close to 0 has an almost uniform distribution of overlaps (with eigenstates). The state away from the energy 0 show a larger concentration of overlap at the edge of the spectrum. The inset in Fig. 12(c) shows the IPR for these states. In the states where $\theta=0.95\pi$, the IPR shows very minor dependence on $\phi$ (similar to what is observed in energy). This explains the absence of a clear QME for this set of states.
Figure 13: Anomalous relaxation in the near-integrable regime at $\tilde{M}=0.8$ for a system with $L=12$. (a) The IPR as a function of the initial state parameters $(\theta,\phi)$. The red and blue vertical lines indicate the two classes of states studied. (b) Relaxation dynamics, $\Delta^{(1)}(t)$, for the nearly-polarized states ($\theta=0.95\pi$). (c) The distribution of overlaps, $|\langle E_{n}|\psi\rangle|^{2}$, with energy eigenstates for several initial states from the $\theta=0.5\pi$ class. (d) Relaxation dynamics for the $\theta=0.5\pi$ class, showing persistent oscillations and multiple crossings. The inset provides a zoomed-in view of the initial relaxation. For all plots, the subsystem size is $L_{A}=4$, and all data are averaged over 500 graph realizations using the Hamiltonian in Eq. 1 with $g=1.0$ and $h=0.1$.
The clear observation of the QME in the chaotic regime is driven by the broad delocalization of initial states across the energy spectrum. A natural question is how this phenomenon behaves as the system approaches the integrable limit at high connectance ($\tilde{M}\to 1$), where the eigenspectrum deviates from chaotic WD statistics. We investigate this interplay at $\tilde{M}=0.8$ (Fig. 13). While the correlation between IPR and energy persists (Fig. 11 and 13(a) ), the overall range of IPR values is smaller, indicating that the initial states are more localized in the eigenbasis as $\tilde{M}\to 1$. Consequently, the nearly-polarized $\theta=0.95\pi$ states still show no strong signature of the QME (Fig. 13(b)). For the $\theta=0.5\pi$ states, however, the dynamics are markedly different from the chaotic case. Instead of a smooth decay, the trace distance exhibits persistent late-time oscillations, as seen in Fig. 13(d). We expect this to be caused by the emergence of approximately conserved charges close to an integrable point. In such near-integrable models, the QME is understood to arise from the coherent propagation of stable quasiparticles, which leads to persistent oscillations and multiple crossings of the relaxation curves [82, 87]. Our results show these characteristic signatures of multiple crossings and oscillatory behavior, and we observe the QME within the timescale studied (Fig. 13(d)).
VII Conclusion and Outlook
In this work, we have investigated the emergence of quantum chaos and complexity in the mixed-field Ising model defined on random Erdős–Rényi graphs, with a particular focus on how the connectance parameter $\tilde{M}$ controls the dynamical and spectral properties of the system. Using a combination of diagnostics, including projected ensemble (PE), partial spectral form factor (pSFF), and Krylov complexity (KC), we provide a detailed characterization of the localized to chaotic to integrable crossover driven by changes in connectance.
The slowdown of deep thermalization at the extreme limits of connectance is signaled by the distance of the PE from the Haar ensemble. In the intermediate chaotic regime, the pSFF exhibits the characteristic ramp-and-plateau behavior indicative of RMT statistics, while the Krylov complexity correspondingly reveals enhanced operator spreading and higher late-time saturation values. Although the studies presented here are limited to modest system sizes, our suite of chaos indicators provides clear and consistent signatures of the crossover from localization to chaos.
We also investigate the relaxation dynamics of initial product states, where initial states further from their equilibrium (“hotter”) thermalize faster than the initial states closer to their equilibrium (“cooler”) under unitary evolution. As $\tilde{M}\to 1$, the signature of the effect is modified, with persistent oscillations causing multiple crossings in the trace distance. Our analysis has focused on the ferromagnetic Ising model. While the dynamical signatures of chaos are expected to be independent of the sign of the couplings, the relaxation dynamics, particularly for specific low-energy initial states, may be altered in the anti-ferromagnetic case, which is a potential direction for future study.
Our results are directly relevant for benchmarking and characterizing thermalization in near-term quantum annealing and simulation platforms. Furthermore, a useful avenue for future work is to focus on the dynamics within the low-energy sector of the spectrum. Since the performance of quantum annealing is sensitive to the localization of low-lying eigenstates [26], applying our suite of chaos diagnostics to this spectral edge could reveal a more direct link between chaos and algorithmic performance.
Acknowledgements.
We thank S Roy for useful discussions and collaborations on an earlier related work. We thank National Supercomputing Mission (NSM) for providing computing resources of “PARAM Brahma” at IISER Pune, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India.
Appendix A Lanczos algorithm
In the main text, we utilized the Lanczos algorithm to generate the basis in Krylov space. The Lanczos algorithm is an iterative method for constructing the orthonormal basis of the Krylov subspace, $\mathcal{K}_{\mathcal{O}}$. Conventional implementation of the algorithm relies on a computationally efficient three-term recurrence relation, where each new vector is made orthogonal only to the two preceding vectors. While theoretically exact, this procedure is numerically unstable when implemented with finite-precision arithmetic, as rounding errors accumulate and lead to a gradual loss of orthogonality among the basis vectors.
To circumvent this issue, we employ the more robust, albeit computationally expensive, method of Lanczos with Full Orthogonalization [88, 68]. In this scheme, each new candidate vector is explicitly made orthogonal to all previously generated basis vectors using a Gram-Schmidt procedure. This ensures a numerically stable orthonormal basis (up to machine precision) at the cost of increased computational complexity, as the number of required inner products scales with the iteration number $n$, rather than being constant. However, for small system sizes, this cost is manageable. The procedure we implement is as follows:
  1. 1.
[…]
    $|\mathcal{O}_{0})=\frac{|\tilde{\mathcal{O}_{0}})}{\sqrt{(\tilde{\mathcal{O}_{0}}|\tilde{\mathcal{O}_{0}})}}$
  2. 2.
    Iteration (for n = 0, 1, 2, …):
    First, generate a new candidate vector by applying the Liouvillian to the most recently generated basis vector:
    $|w_{n+1})=\mathcal{L}|\mathcal{O}_{n})$
    Next, enforce orthogonality. To ensure the new vector is orthogonal to all previously found basis vectors, explicitly subtract the projection onto each of them. For numerical stability, this process is repeated twice (double orthogonalization):
    $|w^{\prime}_{n+1})=|w_{n+1})-\sum_{j=0}^{n}(\mathcal{O}_{j}|w_{n+1})|\mathcal{O}_{j})$
    The next Lanczos coefficient, $b_{n+1}$, is the norm of this fully orthogonalized vector:
    $b_{n+1}=\sqrt{(w^{\prime}_{n+1}|w^{\prime}_{n+1})}$
    If $b_{n+1}$ is smaller than a numerical tolerance (e.g., $10^{-13}$), the Krylov subspace is exhausted, and the algorithm terminates constructing a Krylov subspace of dimension $K$ such that $b_{K}<10^{-13}$. Otherwise, the next orthonormal basis vector is found by normalizing:
    $|\mathcal{O}_{n+1})=\frac{|w^{\prime}_{n+1})}{b_{n+1}}$
This iterative process yields the set of orthonormal basis operators $\{|O_{n}))\}_{n=0}^{K-1}$ and the real, positive Lanczos coefficients $\{b_{n}\}_{n=1}^{K-1}$ that define the tri-diagonal hermitian matrix representation of $\mathcal{L}$.
A key feature of the resulting tridiagonal matrix for $\mathcal{L}$ is that its diagonal elements, $a_{n}=(\mathcal{O}_{n}|\mathcal{L}|\mathcal{O}_{n})$, are identically zero. This arises from the definition of the Liouvillian and the cyclic property of the trace. Assuming the basis operators $|\mathcal{O}_{n}))$ are Hermitian (which is the case if the initial operator is Hermitian), we have:
$a_{n}$ $=(\mathcal{O}_{n}|\mathcal{L}|\mathcal{O}_{n})$
$=\mathrm{Tr}(\mathcal{O}_{n}^{\dagger}[H,\mathcal{O}_{n}])$
$=\mathrm{Tr}(\mathcal{O}_{n}(H\mathcal{O}_{n}-\mathcal{O}_{n}H)$
$=\mathrm{Tr}(\mathcal{O}_{n}H\mathcal{O}_{n})-\mathrm{Tr}(\mathcal{O}_{n}\mathcal{O}_{n}H)=0$
This cancellation ensures that the Liouvillian is purely off-diagonal in the Krylov basis, simplifying the resulting dynamics to a hopping problem on a chain without any on-site potential.
Appendix B Krylov complexity for random single-site operators
In Sec. V, we discussed the Krylov complexity for a simple initial operator (Pauli $Y$). We also saw the effect of the initial size of the operator on its growth. A natural question arises regarding the dependence of the behavior on the choice of operator. Here, we take a random single-site operator as,
$\mathcal{O}_{i}=r_{0}\mathbb{I}+r_{1}X_{i}+r_{2}Y_{i}+r_{3}Z_{i}.$ (22)
Here, $\{r_{0},r_{1},r_{2},r_{3}\}$ are random numbers sampled from a normal distribution. The operator is normalized such that it has a norm of $\sqrt{2}$ (i.e., equal to Pauli operators). In Fig 14, we present the results after averaging over $500$ graphs. For each graph realization, we take a different random initial operator. Averaging is done over both graph realizations and the initial operator. The trend with $\tilde{M}$ holds here as well as shown in Fig 14(a). We compare the asymptotic KC obtained after averaging over random graphs and operators with the scenario where averaging is done only over graphs in Fig. 14(b). The asymptotic KC is smaller for random operators, but the qualitative trend is the same. The time-evolved Krylov wavefunction is shown in Fig. 14(c). The wavefunction shows a clear sign of localization for $\tilde{M}=0.9$. The distribution of Lanczos coefficients is also consistent with the qualitative characteristics observed earlier in Sec. V.
Figure 14: (a) Time evolution of KC and its asymptotic value across multiple $\tilde{M}$. (b) Comparison of asymptotic Krylov complexity as a function of $\tilde{M}$ for random single-site operator and Pauli-$Y$ operator. (c) The asymptotic Krylov wavevector (d) The distribution of Lanczos coefficients. $L=6$ Error bars are shown as the standard error of the mean. All results are presented after averaging over $500$ random graphs and operators.